the probability of guessing the right answer is x by 2 and the probability of guessing the wrong answer is 2 by 3. find the value of x
Answer by stanbon(75887)

the probability of guessing the right answer is x by 2 and the probability of guessing the wrong answer is 2 by 3. find the value of x
-----
Equation:
x/2 = 1-(2/3)
---
x/2 = 1/3
----
x = 2/3
=============
Cheers,
Stan H.
--------------